OpenStudy (anonymous):

A poker hand consists of five cards. a). find the total number of possible five card poker hands. b) a diamond flush is a five card hand consisting of all diamonds. find the number of possible diamond flushes c) find the probability of being dealt a diamond flush

OpenStudy (kropot72):

a) The number of combinations of 52 cards taken 5 at a time is given by: \[\large C(52, 5)=\frac{52!}{5!(52-5)!}=\frac{52\times51\times40\times49\times48}{5\times4\times3\times2\times1}=you\ can\ calculate\]